2016B Day 4: Tipping 2000 Hz

After the secondary was installed yesterday, we ended up going to bed with a bad fiber running to the secondary. So the first order of business was to fix that this morning. All Done.

Katie pulls the ASM cover before installation yesterday.

Laird and Pato installing the calibration return optic (CRO)

Runa and Marco installed the interferometer we will use to calibrate the ASM.

Meanwhile, Enrico installed our new tip-tilt stage, which can go 2000 Hz.

A wild vizcacha came out to watch the sunset.

This is what we watched. No flash.

Take it away Otis:

This entry was posted in Uncategorized. Bookmark the permalink.